OpenStudy (anonymous):

Distribute $3700 among Joe, Tom and Jack so that Joe receives $400 more than Jack and Jack receives twice what Tom receives. Indicate the amount each will receive.

OpenStudy (anonymous):

joe = 400 + 2x jack = 2(x) tom = x x + 2x + 400 + 2x = 3700 5x + 400 = 3700 5x = 3300 x = 660 so tom gets $660 jack gets 2(660) = $1320 and joe gets 400 + 1320 = $1720

OpenStudy (anonymous):

thank you so much . . . you were a great help . . . i will be posting one more story problem . . . whew!